mass of water is 0.3kgAn insulated beaker with negligible mass contains a mass of 0.300 kg of water at a temperature of 84.1 °C. - Part A How many
Answer #1

Let,

  • T_{water}   : be the initial temperature of the water.
  • T_{ice}       : be the initial temperature of the ice.
  • T         : be the final temperature of the system.
  • L          : be the latent heat of fusion of water.
  • C_{ice}     : be the specific heat of ice.
  • C_{water} : be the specific heat of water.
  • m_{water} : mass of the water.
  • m_{ice}     : mass of the ice.

Given:-

  • \small m_{water} = 0.3 \; kg
  • \small T_{water} = 84.1 \: ^{\circ}C
  • \small T_{ice} = -10.3 \: ^{\circ}C
  • \small T = 29.9 \: ^{\circ}C
  • \small C_{water} = 4190 \; J/\left (kg \cdot K \right )
  • \small C_{ice} = 2100 \; J/\left (kg \cdot K \right )
  • \small L = 334 \; kJ/kg = 334 \times 10^{3} \; J/kg

When the system reaches the final temperature T then the total heat flow for the system becomes zero,

Q_{total} = Q_{ice} + Q_{water} = 0

Where,

  • Q_{total}: be the total heat flow for the system.
  • Q_{ice}    : heat for the ice.
  • Q_{water} : heat for water.

Since the ice will go through three stages in this process, hence the Q_{ice} for ice, which has initial temperature of -10.3oC and final temperature 0oC, given by,

Q_{ice} =\left ( m_{ice}\times C_{ice} \times \Delta T_{ice} \right ) + \left ( m_{ice} \times L \right ) + \left ( m_{ice} \times C_{water} \times \Delta T_{melted} \right )

Q_{ice} = m_{ice} \left [C_{ice} \times \Delta T_{ice} + L + C_{water} \times \Delta T_{melted} \right ]

Q_{ice} = m_{ice} \left [2100 \times \left ( 0^{\circ}-\left ( -10.3^{\circ} \right ) \right ) + \left (334 \times 10^{3} \right ) + 4190 \times \left ( 29.9^{\circ} -0 ^{\circ}\right ) \right ]

\mathbf{Q_{ice} = m_{ice} \left (480911 \right )}

This is the heat gain by the ice.

Now, Q_{water} for the water, which drops the temperature from 84.1oC to 29.9oC is give by,

Q_{water} = m_{water} \times C_{water} \times \Delta T_{water}

Q_{water} = 0.3 \times 4190 \times \left ( 29.9^{\circ} - 84.1^{\circ} \right )

\mathbf{Q_{water} = -68129.4 \; J}

The negative sign indicates that the heat is lost by the water.

Now the total heat is given by,

Q_{total} = Q_{ice} + Q_{water} = 0

m_{ice} \left (480911 \right ) -68129.4 = 0

m_{ice} = \frac{68129.4}{480911}

\mathbf{m_{ice} = 0.14 \; kg}

This the value of ice in kilograms which required to drop in to the water to make the final temperature of the system 29.9 oC.